A car slows down from 20m/s with a uniform velocity retardation of 25m/s2 calculate how long it will take the car to reach 5m/s and the distance that will be covered .
1
Expert's answer
2019-11-04T10:11:58-0500

The time is


"t=\frac{V-v}{a}=\frac{20-5}{25}=0.6\ s"

"V^2-v^2=2as"

"20^2-5^2=2(25)s"

The distance that will be covered:


"s=7.5\ m"